OpenStudy (anonymous):

A quiz has twenty questions with seven points awarded for each correct answer, two points deducted for each wrong answer and zero for each question omitted. Jack scores 87 points. How many questions did he omit.

OpenStudy (epoweritheta):

7*CA-2*WA=87; 7*CA=87+2*WA ; note 87+2*wa is a odd number , so see what multiples of 7 are odd and greater than 87 ?? i see 91 , so if WA=2,CA=13 . i also see 105 , WA =9 CA=15 but that is impossible since total qns =20 therefore WA=2,CA=13 .